92 (uC) =1.910 ﹀ ﹀l =-0.750 y (m) 4 91 92 10 x (m) -4 -6 The Vector Nature of the Electrostatic ForceInteractive Exercises 21.06: Practice Using Coulombs Law Consider the default configuration of the two point particles in the simulation (linked below). Notice that neither particle is located at the origin of the coordinate system. In these exercises, we wish to calculate the electrostatic force on particle #1 due to particle #2. Simulation The Vector Nature of the Electrostatic Force Question 1 Many different vectors appear in the simulation. Combine each of the numbered descriptions below with a lettered description for that arrow. Note: the descriptions below are based on the default locations and default charge values for the two particles. 1, 2, 3, 4, 5. 6. 7. Black, open-tipped arrow with tail on particle #1 Blue, open-tipped arrow directed from origin to particle #2 Blue, open-tipped arrow directed from origin to particle #1 Blue, open-tipped arrow directed from particle #2 to particle #1 Purple, open-tipped arrow with tail on particle #1 Purple, closed-tipped arrow directed in-x Purple, closed-tipped arrow directed in -y A. Electrostatic force on particle #1 due to particle #2 B. Position vector for particle #1 relative to origin. C. Unit vector in direction of 12 D. x component of electrostatic force on particle #1 due to particle #2 E. Position vector of particle #1 relative to particle #2 F. Position vector of particle #2 relative to origin. G. y component of electrostatic force on particle #1 due to particle #2. Attempts: 0 of 5 used Question 2 Use information from the simulation to answer these two questions. Note: the questions are based on the default locations and default charge values for the two particles. (a) What is the position of particle #1 relative to position #2 r 12 in unit-vector notation? Fill in the missing numerical values in the following equation r12 2 m (b) What is the magnitude 12 Answer 1: exact number, no tolerance Answer *2: exact number, no tolerance Answer *3: the tolerance is +/-5% Attempts: 0 of 5 usedQuestion 3 Express the unit vector F by filling in the missing numerical values in the following equation aj Answer 1: the tolerance is +/-5% Answer *2: the tolerance is +/-59% Attempts: O of 5 used Question 4 Express the electrostatic force on particle #1 due to particle #2 in unit-vector notation by filling in the missing numerical values (black boxes). Note that units are millinewtons. F 12 2 mN) Answer * 1: the tolerance is +/-5% Answer * 2: the tolerance is +/-5% Attempts: 0 of 5 usedQuestion 7 and Fi2, ? μ what are the x and y components of the electrostatic force on particle #1 due to particle #2,F12 using the sider that controls the sign of the charge of particle #1, set qi 1 mN F12x = F 12.,- Answer * 1: the tolerance is +/-5% 2 mN Answer *2: the tolerance is +/-5%
Answer #1


O-F Q21 r2 =(4,3) - (0,1) = (4,2) 01 - m1 + 2 m . the magnitude Pirl= (4²4? = 4.47m 103 the unit vector ² = [2 = (hê +27) = (

Please, post other parts in another question